# Rational solutions of Painlevé V from Hankel determinants and the asymptotics of their pole locations

### v2

@@
Abstract
In this paper we analyze the asymptotic behaviour of the poles of certain rational solutions of the fifth Painlevé equation. These solutions are constructed by relating the corresponding tau function to a Hankel determinant of a certain sequence of moments. This approach was also used by one of the authors and collaborators in the study of the rational solutions of the second Painlevé equation. More specifically we study the roots of the corresponding polynomial tau function, whose location corresponds to the poles of the associated rational solution. We show that, upon suitable rescaling, the roots fill a well-defined region bounded by analytic arcs when the degree of the polynomial tau function tends to infinity. Moreover we provide an approximate location of these roots within the region in terms of suitable quantization conditions.
Rational solutions of Painlevé V from Hankel determinants
and the asymptotics of their pole locations
M. Balogoun [^1] Malik.Balogoun@concordia.ca M. Bertola [^2] Marco.Bertola@concordia.ca,
Department of Mathematics and Statistics, Concordia University
1455 de Maisonneuve W., Montréal, Québec, Canada H3G 1M8
@@ 1 Introduction
The six Painlevé equations were classified by Painlevé and his student Gambier [53, 29] more than a century ago. They were searching for second order ODEs in the complex plane whose solutions, roughly speaking, have the property that all movable singularities are isolated poles. This property has now become known, and referred to, as the Painlevé property.
While this might have remained a purely mathematical investigation, it was much later recognized that these equations have significant applications in mathematical physics, with the resurgence in the late ’70s with the works connecting with Ising model and conformal field theory [46, 47]. Another momentous resurgence happened in the ’90s when Tracy and Widom [56] used a special second Painlevé transcendent (the Hastings-McLeod solution [31]) to describe the fluctuations of the largest eigenvalue of a large random Hermitean matrix.
Amongst special solutions of the Painlevé equation, the rational ones attract a natural interest; the literature is extensive and seems to start with [59] who discussed rational solutions of the second Painlevé equation and defined a special sequence of polynomials that are now called Vorob’ev-Yablonskii after their discoverers (it appears that Yablonskii defined them slightly earlier but the reference is difficult to find [60]). Rational solutions also appear in semiclassical limits of integrable PDEs; in the one–dimensional sine–Gordon equation near a separatrix, for example, one finds that a suitable scaling of the solution is expressible in terms of a rational solution of the second Painlevé equation [16].
Rational solutions exist for all but the first Painlevé equation. In all five remaining cases II–VI there is a classification of the values of parameters for which there are rational solutions, and for particular cases there are constructions of special families of rational solutions (for the Painlevé II [59, 60, 35], for the Painlevé III, V, VI [45, 42, 57, 50, 51], for Painlevé IV [49]). See a comprehensive overview in [52]. Note that in loc. cit. the conditions for existence of rational solutions of Painlevé VI appears to be only a sufficient condition. However necessary and sufficient conditions exist in [43].
The literature that investigates the asymptotic behaviour of the rational solutions, and the pole distribution thereof, is more recent, probably due to the interest spurred by numerical investigations and the appearance of well-defined patterns; for the zeros of Okamoto polynomials (which are poles of rational solutions of PIV) see [48, 15, 19, 21], for the zeros of Vorobev–Yablonskii polynomials and Painlevé II see [17, 18, 6, 44], for the second Painlevé hierarchy see [4]. There are also studies of explicit families of rational solutions of various Painlevé equations; for Painlevé III [13, 14, 20], for Painleve V [22, 23, 1, 2], for Painlevé IV [15, 21, 40, 41] and for higher Painlevé equations [4, 3].
@@
The isomonodromic approach relies on the general fact that any Painlevé equation appears as the compatibility between a $2\times 2$ system of ODEs with rational coefficients in the complex plane and an additional PDE in an auxiliary parameter [32]. The different solutions are parametrized by (generalized) monodromy data of the ODE, which is the starting point for the Riemann–Hilbert analysis. Typically the degree of the rational solution appears explicitly as one of the parameters in the monodromy data, and can be used as large parameter in the asymptotics. This is the philosophy behind the works [17, 18].
The second approach was used, possibly for the first time, in [6] and then also applied to the generalized Vorob’ev–Yablonski polynomials in [4]. It is also the approach we follow in this paper. The main connection between orthogonal polynomials and equations of Painlevé type was established in [7], where it was shown that Hankel determinants built out of the moments of “semiclassical” moment functionals are always isomonodromic tau functions in the sense of Jimbo-Miwa-Ueno [32]. It was a remark (Rem 5.3 in [7]) that special choices of semiclassical moment functionals lead automatically to tau functions of Painlevé equations (all, except possibly for Painlevé I). In general, however, these solutions correspond to transcendental solutions like, for example, the solutions of PII constructed out of determinants of derivatives of Airy functions, see [34].
@@
Before going into any further detail let us discuss the known literature and results about the rational solutions of the fifth Painlevé equation.
Rational solutions of PV.
The fifth Painlevé equation is the following nonlinear, second order ODE in the complex domain for the unknown function $y(t)$
$y^{\prime\prime}=\left(\frac{1}{2y}+\frac{1}{y-1}\right)(y^{\prime})^{2}-\frac{y^{\prime}}{t}+\frac{(y-1)^{2}}{t^{2}}\left(\alpha y+\frac{\beta}{y}\right)+\frac{\gamma y}{t}+\frac{\delta y(y+1)}{y-1},$ (1.1)
where the prime denotes ${\mathrm{d}}/{{\mathrm{d}}t}$ and $\alpha,\beta,\gamma,\delta\in{\mathbb{C}}$ are parameters: we shall refer to (1.1) as $P_{5}(\alpha,\beta,\gamma,\delta)$. The equation admits certain symmetries that change the value of the parameters; if $y(t)$ is a solution of $P_{5}(\alpha,\beta,\gamma,\delta)$ then
  1. 1.
    $y(-t)$ is a solution of $P_{5}(\alpha,\beta,-\gamma,\delta)$.
  2. 2.
    $\frac{1}{y(t)}$ is a solution of $P_{5}(-\beta,-\alpha,-\gamma,\delta)$.
  3. 3.
    $y(\lambda t)$ is a solution of $P_{5}(\alpha,\beta,\lambda\gamma,\lambda^{2}\delta)$, for any $\lambda\in{\mathbb{C}}\setminus\{0\}$.
Using the above symmetries the analysis is reduced to only two families; the family where $\delta=0$ (which is called “degenerate” and can be reduced to Painlevé III) and the case $\delta\neq 0$ which, by virtue of the last of the above symmetries, is customarily set to $\delta=-\frac{1}{2}$.
The classification of rational solutions is contained in [36] where the authors show that rational solutions exist only if the parameters satisfy certain relations. More precisely (paraphrasing and condensing their results)
Theorem 1.1 ([36], Thm. 1.1, Thm. 1.2)
The equation $P_{5}(\alpha,\beta,\gamma,-\frac{1}{2})$ (1.1) admits rational solutions if and only if there are integers $k,m\in{\mathbb{Z}}$ such that
@@
    $\alpha=\frac{1}{2}(\gamma+k)^{2};\ \ \beta=-\frac{m^{2}}{2}$, $k+m$ odd, and $\alpha\neq 0$ when $|k|<m$.
@@
    $\alpha=\frac{1}{2}m^{2};\ \ \beta=-\frac{(\gamma+k)^{2}}{2}$, $k+m$ odd, and $\beta\neq 0$ when $|k|<m$.
@@
    $\beta=-\frac{1}{2}(\alpha_{1}+m)^{2},\gamma=k$ with $\alpha_{1}^{2}=2\alpha$ so that $m\geq 0$ and $k+m$ even.
@@
In cases (I) and (II) the solution is unique if $\gamma\not\in{\mathbb{Z}}$ and there are at most two rational solutions otherwise.
Observe that the cases (I) and (II) of the above result are really the same family up to the application of the symmetry $y(t)\leftrightarrow\frac{1}{y(-t)}$ which transforms solution of $P_{5}(\alpha,\beta,\gamma,-\frac{1}{2})$ into solutions of $P_{5}(-\beta,-\alpha,\gamma,-\frac{1}{2})$.
In the recent [23] the authors construct the rational solutions corresponding to the case (II) of [36] above (which is case (i) in Thm. 4.1 of [23]). More precisely they construct the tau functions
$\tau^{(\mu)}_{m,n}:=\det\left[\left(t\frac{{\mathrm{d}}}{{\mathrm{d}}t}\right)^{j+k}L^{(n+\mu)}_{m+n}(t)\right]_{j,k=0}^{n-1}$ (1.2)
@@
$L^{(\alpha)}_{n}\left(t\right)=\sum_{\ell=0}^{n}\frac{{\left(\alpha+\ell+1\right)_{n-\ell}}}{(n-\ell)!\;\ell!}(-x)^{\ell}=\frac{{\left(\alpha+1\right)_{n}}}{n!}{{}_{1}F_{1}}\left({-n\atop\alpha+1};t\right).$ (1.3)
@@
$P_{5}\left(\frac{m^{2}}{2},-\frac{(m+2n+1+\mu)^{2}}{2},\mu,-\frac{1}{2}\right),$ (1.4)
which is the case (II) in Thm. 1.1. Many interesting properties are discussed but no description of the asymptotic behaviour for large degree is undertaken.
Results.
In the present paper we will provide a description for the tau function of the case (I) of Thm. 1.1, with the main goal of describing the distribution of its zeroes as $n\to\infty$. More precisely, we construct the tau function corresponding to the parameters
$k=-3n-{K}-1,\ \ m=n+{K},\ \ \gamma=1+2n+{K}-{\rho},\ \ \ n\in{\mathbb{N}},\ \ {K}\in{\mathbb{Z}},\ \ {\rho}\in{\mathbb{C}},$ (1.5)
namely
$\alpha=\frac{(n+{{\rho}})^{2}}{2};\ \ \beta=-\frac{(n+{{K}})^{2}}{2};\ \ \gamma=1+2n+{{K}}-{{\rho}}.$ (1.6)
@@ Results.
$P_{5}\left(\frac{(n+{{\rho}})^{2}}{2},-\frac{(n+{{K}})^{2}}{2},1+2n+{{K}}-{{\rho}},-\frac{1}{2}\right).$ (1.7)
Remark 1.1
The rational solutions we consider here are for case (I) while those in [23] are for case (II) of Thm. 1.1. Of course the two cases are related: at the level of the transcendents the solutions are related by the simple transformation $y(t)\mapsto\frac{1}{y(t)}$ and $(\alpha,\beta,\gamma)\mapsto(-\beta,-\alpha,-\gamma)$, but at the level of the tau functions the relation is slightly less straightforward and given in Rem. 2.1 below. The comparison of the parameters $\mu,n,m$ of [23] (which we now have to rename $\widetilde{\mu},\widetilde{n},\widetilde{m}$ to avoid confusion) is then as follows
$\widetilde{m}=n+{{K}},\ \ \widetilde{n}=n,\ \ \ \widetilde{\mu}=-1-2n-{{K}}+{{\rho}}.$ (1.8)
$\triangle$
To construct this $\tau$-function we consider the following sequence of moments depending on the parameters ${K}\in{\mathbb{Z}}$, ${\rho}\in{\mathbb{C}}$
$\mu_{j}(t):=\oint_{|z|=R}z^{j}z^{{K}}\left(1-\frac{1}{z}\right)^{{\rho}}{\rm e}^{\frac{t}{z}}{\mathrm{d}}z,\ \ \ \ \ j=0,1,2,\dots.$ (1.9)
Here $R>1$ can be chosen arbitrarily, thanks to Cauchy’s residue theorem. We interpret the $\mu_{j}$’s as the moments of the complex measure ${\mathrm{d}}\mu(z)=z^{{K}}\left(1-\frac{1}{z}\right)^{{\rho}}{\rm e}^{\frac{t}{z}}{\mathrm{d}}z$ on $|z|=R$. Observe that the $\mu_{j}(t)$’s are polynomials in $t$, and can also be obtained from a generating function
$\sum_{j=0}^{\infty}\mu_{j}(t)\frac{s^{j}}{j!}=\oint_{|z|=R}z^{{K}}\left(1-\frac{1}{z}\right)^{{\rho}}{\rm e}^{\frac{t}{z}+sz}{\mathrm{d}}z$ (1.10)
@@
$F(s;t):=-2i\pi s^{-{K}-1}\left(1-s\right)^{{\rho}}{\rm e}^{st}=\sum_{j=0}^{\infty}\mu_{j}(t)\frac{s^{j}}{j!}.$ (1.11)
@@
$\tau_{n}(t;{K},{\rho}):=\det\left[\mu_{a+b-2}(t)\right]_{a,b=1}^{n},$ (1.12)
is a tau function of an equation of Painlevé type. The matching with PV is explained in Section 2.2. More explicitly, from the identification it will follow that it satisfies the $\sigma$–form of the PV equation ([33], App. C) in the following way: let us define the quantities
$H_{{}_{V}}$ $:=\frac{{\mathrm{d}}}{{\mathrm{d}}t}\ln\tau_{n}(t;{K},{\rho})+\frac{\rho}{2}$ (1.13)
$\sigma(t)$ $:=tH_{V}+\frac{t}{2}(\theta_{0}+\theta_{\infty})+\frac{(\theta_{0}+\theta_{\infty})^{2}-\theta_{1}^{2}}{4}$ (1.14)
$\theta_{0}$ $:=2n+{{K}};\qquad\theta_{1}:=-{{\rho}};\qquad\theta_{\infty}:={{\rho}}-{{K}}.$ (1.15)
Then the $\sigma$–form of PV is the following ODE for $\sigma$ defined in (1.14)
$\left(t\frac{{\mathrm{d}}^{2}\sigma}{{\mathrm{d}}t^{2}}\right)^{2}=$ $\left(\sigma-t\frac{{\mathrm{d}}\sigma}{{\mathrm{d}}t}+2\left(\frac{{\mathrm{d}}\sigma}{{\mathrm{d}}t}\right)^{2}-(\theta_{\infty}+2\theta_{0})\frac{{\mathrm{d}}\sigma}{{\mathrm{d}}t}\right)^{2}+$
$-4\left(\frac{{\mathrm{d}}\sigma}{{\mathrm{d}}t}\right)\left(\frac{{\mathrm{d}}\sigma}{{\mathrm{d}}t}-\frac{\theta_{0}-\theta_{1}+\theta_{\infty}}{2}\right)\left(\frac{{\mathrm{d}}\sigma}{{\mathrm{d}}t}-\theta_{0}\right)\left(\frac{{\mathrm{d}}\sigma}{{\mathrm{d}}t}-\frac{\theta_{0}+\theta_{1}+\theta_{\infty}}{2}\right).$ (1.16)
In the second part of the paper we exploit the connection between $\tau_{n}(t)$ and Hankel determinants to describe the asymptotic location of its zeros as $n\to\infty$.
More precisely we describe the asymptotic behaviour of the zeros of the polynomials
$\mathcal{T}_{n}(s):=\tau_{n}(ns;{K},{\rho}),\ \ \ {K}\in{\mathbb{Z}},\ \ {\rho}\not\in{\mathbb{Z}},$ (1.17)
as $n$ tends to infinity in the $s$–plane (these zeros are the same as the zeros of $\tau_{n}$ but homothetically rescaled by a factor $1/n$). We observe that we do not consider a double–scaling limit where the either one of the parameters ${K}$, ${\rho}$ or both are proportional to $n$ as $n\to\infty$. While the general setup is conducive to such an analysis, the details of the construction of the $g$–function would have to be changed significantly and this is deferred to a future publication. The logic of the analysis is parallel to the one employed in [6, 4] and the main goal is to explain the emerging shape which clearly appears, see Fig. 1.
@@
$n=16,\ {\rho}=\frac{3}{100}+\frac{13i}{100}$
$n=16,\ {\rho}=\frac{101}{100}+\frac{33i}{100}$

### v3

@@
[…]
Rational Solutions of Painlevé V
[…]
Rational Solutions of Painlevé V
from Hankel Determinants and the Asymptotics
of Their Pole Locations
[…]
Malik BALOGOUN and Marco BERTOLA
[…]
M. Balogoun and M. Bertola
[…]
Department of Mathematics and Statistics, Concordia University,
[…]
Received November 14, 2024, in final form October 28, 2025; Published online November 14, 2025
[…]
In this paper, we analyze the asymptotic behaviour of the poles of certain rational solutions of the fifth Painlevé equation. These solutions are constructed by relating the corresponding tau function to a Hankel determinant of a certain sequence of moments. This approach was also used by one of the authors and collaborators in the study of the rational solutions of the second Painlevé equation. More specifically, we study the roots of the corresponding polynomial tau function, whose location corresponds to the poles of the associated rational solution. We show that, upon suitable rescaling, the roots asymptotically fill a region bounded by analytic arcs when the degree of the polynomial tau function tends to infinity and the other parameters are kept fixed. Moreover, we provide an approximate location of these roots within the region in terms of suitable quantization conditions.
[…]
Painlevé equations; rational solutions; asymptotic analysis; Riemann–Hilbert problems
[…]
33E17; 34M55; 33C47
@@ 1 Introduction
The six Painlevé equations were classified by Painlevé and his student Gambier [30, 54] more than a century ago. They were searching for second-order ODEs in the complex plane whose solutions, roughly speaking, have the property that all movable singularities are isolated poles. This property has now become known, and referred to, as the Painlevé property.
While this might have remained a purely mathematical investigation, it was much later recognized that these equations have significant applications in mathematical physics, with the resurgence in the late 70s with the works connecting with Ising model and conformal field theory [45, 46]. Another momentous resurgence happened in the 90s when Tracy and Widom [57] used a special second Painlevé transcendent (the Hastings–McLeod solution [31]) to describe the fluctuations of the largest eigenvalue of a large random Hermitean matrix.
Amongst special solutions of the Painlevé equation, the rational ones attract a natural interest. The literature is extensive and seems to start with [60] who discussed rational solutions of the second Painlevé equation and defined a special sequence of polynomials that are now called Vorob’ev–Yablonskii after their discoverers (it appears that Yablonskii defined them slightly earlier but the reference is difficult to find [61]). Rational solutions also appear in semiclassical limits of integrable PDEs. In the one-dimensional sine-Gordon equation near a separatrix, for example, one finds that a suitable scaling of the solution is expressible in terms of a rational solution of the second Painlevé equation [16].
Rational solutions exist for all but the first Painlevé equation. In all five remaining cases II–VI, there is a classification of the values of parameters for which there are rational solutions, and for particular cases there are constructions of special families of rational solutions (for Painlevé II [35, 60, 61], for Painlevé III, V, VI [42, 43, 48, 51, 52, 58], for Painlevé IV [50]). See a comprehensive overview in [53] (https://dlmf.nist.gov/32.8). Note that in loc. cit. the conditions for existence of rational solutions of Painlevé VI appears to be only a sufficient condition. However, necessary and sufficient conditions exist in [44].
The literature that investigates the asymptotic behaviour of the rational solutions, and the pole distribution thereof, is more recent, probably due to the interest spurred by numerical investigations and the appearance of well-defined patterns. For the zeros of Okamoto polynomials (which are poles of rational solutions of PIV) see [15, 19, 20, 49], for the zeros of Vorob’ev–Yablonskii polynomials and Painlevé II see [6, 17, 18, 24, 47], for the second Painlevé hierarchy see [4, 24]. There are also studies of explicit families of rational solutions of various Painlevé equations: for Painlevé III [13, 14, 21], for Painlevé V [1, 3, 22, 23], for Painlevé IV [15, 20, 40, 41] and for higher Painlevé equations [2, 4].
@@ 2. $(2)$
The isomonodromic approach relies on the general fact that any Painlevé equation appears as the compatibility between a $2\times 2$ system of ODEs with rational coefficients in the complex plane and an additional PDE in an auxiliary parameter [32]. The different solutions are parametrized by (generalized) monodromy data of the ODE, which is the starting point for the Riemann–Hilbert analysis. Typically, the degree of the rational solution appears explicitly as one of the parameters in the monodromy data, and can be used as large parameter in the asymptotics. This is the philosophy behind the works [17, 18].
The second approach was used, possibly for the first time, in [6] and then also applied to the generalized Vorob’ev–Yablonskii polynomials in [4]. It is also the approach we follow in this paper. The main connection between orthogonal polynomials and equations of Painlevé type was established in [7], where it was shown that Hankel determinants built out of the moments of “semiclassical” moment functionals are always isomonodromic tau functions in the sense of Jimbo–Miwa–Ueno [32]. It was a remark [7, Remark 5.3] that special choices of semiclassical moment functionals lead automatically to tau functions of Painlevé equations (all, except possibly for Painlevé I). In general, however, these solutions correspond to transcendental solutions like, for example, the solutions of PII constructed out of determinants of derivatives of Airy functions, see [34].
@@
Before going into any further detail, let us discuss the known literature and results about the rational solutions of the fifth Painlevé equation.
Rational solutions of PV. The fifth Painlevé equation is the following nonlinear, second-order ODE in the complex domain for the unknown function $y(t)$
$y^{\prime\prime}=\left(\frac{1}{2y}+\frac{1}{y-1}\right)\bigl(y^{\prime}\bigr)^{2}-\frac{y^{\prime}}{t}+\frac{(y-1)^{2}}{t^{2}}\left(\alpha y+\frac{\beta}{y}\right)+\frac{\gamma y}{t}+\frac{\delta y(y+1)}{y-1},$ (1.1)
where the prime denotes ${\mathrm{d}}/{{\mathrm{d}}t}$ and $\alpha,\beta,\gamma,\delta\in{\mathbb{C}}$ are parameters: we shall refer to (1.1) as $P_{5}(\alpha,\beta,\gamma,\delta)$. The equation admits certain symmetries that change the value of the parameters. If $y(t)$ is a solution of $P_{5}(\alpha,\beta,\gamma,\delta)$, then
  1. (1)
    y(t)y(-t) is a solution of $P_{5}(\alpha,\beta,-\gamma,\delta)$,
  2. (2)
    $\frac{1}{y(t)}$ is a solution of $P_{5}(-\beta,-\alpha,-\gamma,\delta)$,
  3. (3)
    $y(\lambda t)$ is a solution of $P_{5}\bigl(\alpha,\beta,\lambda\gamma,\lambda^{2}\delta\bigr)$, for any $\lambda\in{\mathbb{C}}\setminus\{0\}$.
Using the above symmetries, the analysis is reduced to only two families: the family where $\delta=0$ (which is called “degenerate” and can be reduced to Painlevé III) and the case $\delta\neq 0$ which, by virtue of the last of the above symmetries, is customarily set to $\delta=-\frac{1}{2}$.
The classification of rational solutions is contained in [36], where the authors show that rational solutions exist only if the parameters satisfy certain relations. More precisely (paraphrasing and condensing their results), we have the following.
Theorem 1.1 ([36, Theorems 1.1 and 1.2]).
The equation $P_{5}\bigl(\alpha,\beta,\gamma,-\frac{1}{2}\bigr)$ (1.1) admits rational solutions if and only if there are integers $k,m\in{\mathbb{Z}}$ such that
@@ 1. $(I)$
    $\alpha=\frac{1}{2}(\gamma+k)^{2}$, $\beta=-\frac{m^{2}}{2}$, $k+m$ odd, and $\alpha\neq 0$ when $|k|<m$,
@@ 2. $(II)$
    $\alpha=\frac{1}{2}m^{2}$, $\beta=-\frac{(\gamma+k)^{2}}{2}$, $k+m$ odd, and $\beta\neq 0$ when $|k|<m$,
@@ 3. $(III)$
    $\beta=-\frac{1}{2}(\alpha_{1}+m)^{2}$, $\gamma=k$ with $\alpha_{1}^{2}=2\alpha$ so that $m\geq 0$ and $k+m$ even,
@@ 4. $(IV)$
In cases $(I)$ and $(II)$, the solution is unique if $\gamma\not\in{\mathbb{Z}}$ and there are at most two rational solutions otherwise.
Observe that the cases (I) and (II) of the above result are really the same family up to the application of the symmetry $y(t)\leftrightarrow\frac{1}{y(-t)}$ which transforms solution of $P_{5}\bigl(\alpha,\beta,\gamma,-\frac{1}{2}\bigr)$ into solutions of $P_{5}\bigl(-\beta,-\alpha,\gamma,-\frac{1}{2}\bigr)$.
In the recent [23], the authors construct the rational solutions corresponding to the case (II) of [36] above (which is case (i) in [23, Theorem 4.1]). More precisely, they construct the tau functions
$\tau^{(\mu)}_{m,n}:=\det\left[\left(t\frac{{\mathrm{d}}}{{\mathrm{d}}t}\right)^{j+k}L^{(n+\mu)}_{m+n}(t)\right]_{j,k=0}^{n-1},$ (1.2)
@@
$L^{(\alpha)}_{n}\left(t\right)=\sum_{\ell=0}^{n}\frac{{\left(\alpha+\ell+1\right)_{n-\ell}}}{(n-\ell)!\;\ell!}(-x)^{\ell}=\frac{{\left(\alpha+1\right)_{n}}}{n!}{{}_{1}F_{1}}\left({-n\atop\alpha+1};t\right).$
@@
$P_{5}\left(\frac{m^{2}}{2},-\frac{(m+2n+1+\mu)^{2}}{2},\mu,-\frac{1}{2}\right),$
which is the case (II) in Theorem 1.1. Many interesting properties are discussed but no description of the asymptotic behaviour for large degree is undertaken.
Results. In the present paper, we will provide a description for the tau function of the case (I) of Theorem 1.1, with the main goal of describing the distribution of its zeros as $n\to\infty$. More precisely, we construct the tau function corresponding to the parameters k=3nK1k=-3n-{K}-1, m=n+Km=n+{K}, $\gamma=1+2n+{K}-{\rho}$, $n\in{\mathbb{N}}$, ${K}\in{\mathbb{Z}}$, $\rho$$\in{\mathbb{C}}$, namely
$\alpha=\frac{(n+{{\rho}})^{2}}{2},\qquad\beta=-\frac{(n+{{K}})^{2}}{2},\qquad\gamma=1+2n+{{K}}-{{\rho}}.$
@@
$P_{5}\left(\frac{(n+{{\rho}})^{2}}{2},-\frac{(n+{{K}})^{2}}{2},1+2n+{{K}}-{{\rho}},-\frac{1}{2}\right).$
Remark 1.2.
The rational solutions we consider here are for case (I) while those in [23] are for case (II) of Theorem 1.1. Of course, the two cases are related: at the level of the transcendents the solutions are related by the simple transformation $y(t)\mapsto\frac{1}{y(t)}$ and $(\alpha,\beta,\gamma)\mapsto(-\beta,-\alpha,-\gamma)$, but at the level of the tau functions the relation is slightly less straightforward and given in Remark 2.2 below. The comparison of the parameters $\mu$, $n$, $m$ of [23] (which we now have to rename $\widetilde{\mu}$, $\widetilde{n}$, $\widetilde{m}$ to avoid confusion) is then as follows $\widetilde{m}=n+{{K}}$, $\widetilde{n}=n$, $\widetilde{\mu}=-1-2n-{{K}}+{{\rho}}$.
To construct this $\tau$-function, we consider the following sequence of moments depending on the parameters ${K}\in{\mathbb{Z}}$, $\rho$$\in{\mathbb{C}}$:
$\mu_{j}(t):=\oint_{|z|=R}z^{j}z^{{K}}\left(1-\frac{1}{z}\right)^{{\rho}}{\rm e}^{\frac{t}{z}}{\mathrm{d}}z,\qquad j=0,1,2,\dots.$ (1.3)
As we detail in Remark 2.2, the moments $\mu_{j}(t)$ are Laguerre polynomials up to a sign. Here $R>1$ can be chosen arbitrarily, thanks to Cauchy’s residue theorem. We interpret the $\mu_{j}$’s as the moments of the complex measure ${\mathrm{d}}\mu(z)=z^{{K}}\left(1-\frac{1}{z}\right)^{{\rho}}{\rm e}^{\frac{t}{z}}{\mathrm{d}}z$ on $|z|=R$. Observe that the $\mu_{j}(t)$’s are polynomials in $t$, and can also be obtained from a generating function
$\sum_{j=0}^{\infty}\mu_{j}(t)\frac{s^{j}}{j!}=\oint_{|z|=R}z^{{K}}\left(1-\frac{1}{z}\right)^{{\rho}}{\rm e}^{\frac{t}{z}+sz}{\mathrm{d}}z$
@@
$F(s;t):=-2{\rm i}\pi s^{-{K}-1}\left(1-s\right)^{{\rho}}{\rm e}^{st}=\sum_{j=0}^{\infty}\mu_{j}(t)\frac{s^{j}}{j!}.$
@@
$\tau_{n}(t;{K},{\rho}):=\det\left[\mu_{a+b-2}(t)\right]_{a,b=1}^{n}$
is a tau function of an equation of Painlevé type. The matching with PV is explained in Section 2.2. More explicitly, from the identification it will follow that it satisfies the $\sigma$-form of the PV equation [33, Appendix C] in the following way. Let us define the quantities
$H_{{}_{V}}:=\frac{{\mathrm{d}}}{{\mathrm{d}}t}\ln\tau_{n}(t;{K},{\rho})+\frac{\rho}{2},\qquad\sigma(t):=tH_{V}+\frac{t}{2}(\theta_{0}+\theta_{\infty})+\frac{(\theta_{0}+\theta_{\infty})^{2}-\theta_{1}^{2}}{4},$
$\theta_{0}:=2n+{{K}},\qquad\theta_{1}:=-{{\rho}},\qquad\theta_{\infty}:={{\rho}}-{{K}}.$ (1.4)
Then the $\sigma$-form of PV is the following ODE for $\sigma$ defined in (1.4)
$\left(t\frac{{\mathrm{d}}^{2}\sigma}{{\mathrm{d}}t^{2}}\right)^{2}={}$ $\left(\sigma-t\frac{{\mathrm{d}}\sigma}{{\mathrm{d}}t}+2\left(\frac{{\mathrm{d}}\sigma}{{\mathrm{d}}t}\right)^{2}-(\theta_{\infty}+2\theta_{0})\frac{{\mathrm{d}}\sigma}{{\mathrm{d}}t}\right)^{2}$
$-4\left(\frac{{\mathrm{d}}\sigma}{{\mathrm{d}}t}\right)\left(\frac{{\mathrm{d}}\sigma}{{\mathrm{d}}t}-\frac{\theta_{0}-\theta_{1}+\theta_{\infty}}{2}\right)\left(\frac{{\mathrm{d}}\sigma}{{\mathrm{d}}t}-\theta_{0}\right)\left(\frac{{\mathrm{d}}\sigma}{{\mathrm{d}}t}-\frac{\theta_{0}+\theta_{1}+\theta_{\infty}}{2}\right).$
In the second part of the paper, we exploit the connection between $\tau_{n}(t)$ and Hankel determinants to describe the asymptotic location of its zeros as $n\to\infty$.
More precisely, we describe the asymptotic behaviour of the zeros of the polynomials
$\mathcal{T}_{n}(s):=\tau_{n}(ns;{K},{\rho}),\qquad{K}\in{\mathbb{Z}},\quad{\rho}\not\in{\mathbb{Z}},$ (1.5)
as $n$ tends to infinity in the $s$-plane (these zeros are the same as the zeros of $\tau_{n}$ but homothetically rescaled by a factor $1/n$). We observe that we do not consider a double-scaling limit where the either one of the parameters K{K}, $\rho$} or both are proportional to $n$ as $n\to\infty$. While the general setup is conducive to such an analysis, the details of the construction of the $g$-function would have to be changed significantly and this is deferred to a future publication. The logic of the analysis is parallel to the one employed in [4, 6] and the main goal is to explain the emerging shape which clearly appears, see Figure 1.
@@
$n=16,\ {\rho}=\frac{3}{100}+\frac{13{\rm i}}{100}$
$n=16,\ {\rho}=\frac{101}{100}+\frac{33{\rm i}}{100}$